OpenStudy (anonymous):

hot to convert the equation -x^2+3=(4x-1)^2 into ax^2+bx+c=0 form

OpenStudy (lgbasallote):

first step: simplify (4x-1)^2

OpenStudy (anonymous):

k so it would be 1 or 16x-1?

OpenStudy (anonymous):

ugh im so confused:l

OpenStudy (lgbasallote):

are you talking about the answer?

OpenStudy (anonymous):

yeah, i wanna put the equation into standard form but i have no idea how

OpenStudy (lgbasallote):

well like i said you have to expand (4x-1)^2 first...what is the expansion of this?

OpenStudy (anonymous):

16x-1?

OpenStudy (anonymous):

i dont know what to do after you get -x^2+3=16x-1

OpenStudy (lgbasallote):

it's not 16x -1

OpenStudy (anonymous):

aah okay, so what is it?

OpenStudy (lgbasallote):

(x+1)^2 = (x+1)(x+1) = x^2 + x + x + 1 = x^2 + 2x + 1 (x+2)^2 = (x+2)(x+2) = x^2 + 2x + 2x + 4 = x^2 + 4x + 4 (x-3)^2 = (x-3)(x-3) = x^2 - 3x - 3x + 9 = x^2 - 6x + 9 so what does (4x-1)^2 equal to?

OpenStudy (anonymous):

i have no idea how you even did any of those equations

OpenStudy (anonymous):

r u sure this is how you put the equation into ax^2+bx+c+0?

OpenStudy (lgbasallote):

yep

OpenStudy (lgbasallote):

\[\huge (4x-1)^2 \implies (4x-1)(4x-1)\] now use FOIL on (4x-1)(4x-1)

OpenStudy (anonymous):

all ik from an example is 0x^2+3=(2x-1)^2 converted is 5x^2-4x-2=0

OpenStudy (lgbasallote):

did you write something wrong there?

OpenStudy (anonymous):

so how is that supposed to help me?

OpenStudy (anonymous):

insted of a 0 it was supposed to be - sorry

OpenStudy (lgbasallote):

hmm okay...how about i show you how they got that...maybe you'll get an idea

OpenStudy (anonymous):

ok, thankyou sooooo much!

OpenStudy (anonymous):

i know how they got eveything in the equation except the 5x^2 part

OpenStudy (lgbasallote):

\[-x^2 +3 = (2x-1)^2\] expand (2x-1)^2 first \[(2x-1)^2 \implies (2x-1)(2x-1) \implies 4x^2 - 2x - 2x + 1 \implies 4x^2 - 4x + 1\] so the equation a while ago becomes \[-x^2 + 3 = 4x^2 - 4x + 1\] add x^2 to both sides \[3 = 5x^2 - 4x + 1\] now subtract 3 from both sides \[0 = 5x^2 - 4x - 2\]

OpenStudy (lgbasallote):

does that help?

OpenStudy (anonymous):

ummm one sec

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

so the answer to -x^2+3=(4x-1)^2 is 17x^2-8x-3=0?

OpenStudy (lgbasallote):

very close...except for -3...recheck your work

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

17x^2-8x-2???

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

YES! thank you so much:)

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (anonymous):

caaaann i ask another question?

OpenStudy (lgbasallote):

sure

OpenStudy (anonymous):

how would you find the solutions to the equation 2x^2+5x+8=5?

OpenStudy (anonymous):

would you convert it to 2x^2+5x+3=0 first?

OpenStudy (lgbasallote):

yep. then you factor if possible

OpenStudy (anonymous):

ok sooo how do you do that? haha u probs think im hella stupid:P

OpenStudy (lgbasallote):

nahh

OpenStudy (lgbasallote):

if you're not used to factoring you can just use quadratic formula \[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] do you know how to use it?

OpenStudy (anonymous):

yeah

OpenStudy (lgbasallote):

great. so use it. then tell me what you get

OpenStudy (anonymous):

okay soooo -3 and 2?

OpenStudy (anonymous):

haha im thinkin its wrong

OpenStudy (lgbasallote):

hmm...yeah..it is...

OpenStudy (anonymous):

ooh wait, sorry wrong problem i actually need to do x^2+20=-9x

OpenStudy (lgbasallote):

hint: a = 2 b= 5 c = 3 \[x = \frac{-5 \pm \sqrt{5)^2 - 4(2)(3)}}{2(2)}\]

OpenStudy (lgbasallote):

lol

OpenStudy (lgbasallote):

find the solutions?

OpenStudy (anonymous):

so it would be -4 and -5?

OpenStudy (anonymous):

haha yeah im not the brightest at math:)

OpenStudy (lgbasallote):

yup -4 and -5 x^2 + 9x + 20 = 0 (x+5)(x+4) = 0 so roots are x = -5 and x = -4

OpenStudy (anonymous):

so insted of doing the super long way you can just find the lowest factors of 20?

OpenStudy (lgbasallote):

haha yup...but expressions arent always facorable...and finding the factors isnt that easy either

OpenStudy (anonymous):

darn:l

OpenStudy (anonymous):

are you like a teacher or somthin?

OpenStudy (lgbasallote):

close to that haha

OpenStudy (anonymous):

haha close?

OpenStudy (lgbasallote):

yeah..something like that =))

OpenStudy (anonymous):

lol gotcha

OpenStudy (anonymous):

another question?

OpenStudy (lgbasallote):

go for it

OpenStudy (anonymous):

what number would you add to complete the square for the equation x^2+6x=0

OpenStudy (lgbasallote):

do you know how to complete the square?

OpenStudy (anonymous):

no idea haha

OpenStudy (lgbasallote):

hmm that's okay...can you identify what the COEFFICENT of x is?

OpenStudy (anonymous):

1?

OpenStudy (lgbasallote):

1 is the coefficient of x^2...im asking he coefficient of x

OpenStudy (anonymous):

ooh 6?

OpenStudy (lgbasallote):

right. now divide it by 2

OpenStudy (anonymous):

3

OpenStudy (lgbasallote):

now square it

OpenStudy (anonymous):

12

OpenStudy (anonymous):

nooo

OpenStudy (lgbasallote):

hehe

OpenStudy (anonymous):

9

OpenStudy (anonymous):

wait is it seriously that easy????

OpenStudy (anonymous):

haha yea im a weird one:)

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

wow

OpenStudy (lgbasallote):

i mean yes it is seriously that easy...not that your a weird one

OpenStudy (anonymous):

so then how would you do x^2+12x=9?

OpenStudy (lgbasallote):

how would you?

OpenStudy (anonymous):

haha nah i didnt say wow cuz of the yup i ment wow cuz it was that easy

OpenStudy (anonymous):

to complete the square

OpenStudy (lgbasallote):

oh...the whole process?

OpenStudy (anonymous):

yes pleasseee

OpenStudy (lgbasallote):

so what will you add to both sides?

OpenStudy (anonymous):

add?

OpenStudy (lgbasallote):

like the question earlier

OpenStudy (lgbasallote):

the question a while ago was what number would you add to complete the square for the equation x^2+6x=0 this time it's what number would you add to complete the square for the equation x^2+12x = 9

OpenStudy (anonymous):

ooh so it would be 36?

OpenStudy (lgbasallote):

yup so add 36 to BOTH sides x^2 + 12x + 36 = 9 + 36 simplify it

OpenStudy (anonymous):

x^2+12x-9=0?

OpenStudy (anonymous):

how would you complete the square for 2x^2-28x-2=-50 and then solve for x?